OpenStudy (airyana1114):

Let f(x)=x^2-16. Find f^-1(x). Can someone check my work?

OpenStudy (airyana1114):

\[f(x)=x^{2}-16 \] Step1: Replace f(x) with y. \[y=x ^{2}-16\] Step 2: Switch my x and y variables. \[x=y ^{2}-16\] Step 3: Add 16 to both sides of the equation. \[x+16=y ^{2}\] Step 4: Take the square root of both sides of the equation. \[\sqrt{x+16}=y\] Step 5: Replace y with f(x). \[\sqrt{x+16}=f ^{-1}(x)\] Did I do this right?

OpenStudy (johnweldon1993):

Correct indeed :)

OpenStudy (airyana1114):

Awesome, thank you! @johnweldon1993

OpenStudy (johnweldon1993):

No problem great work!